OpenStudy (anonymous):

In Algebra can you help me solve y=1/2x Slope and y-Inter.

OpenStudy (anonymous):

slope is 1/2 y-intercept is 0

jimthompson5910 (jim_thompson5910):

y=1/2x really looks like y=1/2x + 0 and it's in the form y = mx+b where m is the slope and (0, b) is the y-intercept Since m = 1/2 and b = 0, the slope is 1/2 and the y-intercept is (0, 0)

OpenStudy (saifoo.khan):

My answer is same as moneybird's.

OpenStudy (anonymous):

My vote for jim...nice work jim

OpenStudy (anonymous):

my answer is same as saifoo